# On the first test of the Weak Equivalence Principle in low Earth orbit

### v4

@@ Abstract
The Weak Equivalence Principle is the founding pillar of General Relativity and as such should be verified as precisely as possible. The Microscope experiment tested it in low-Earth orbit, finding that Pt and Ti test masses fall toward Earth with the same acceleration to about 1 part in $10^{15}$, an improvement of about two orders of magnitude over ground tests. Space missions, even if small, are expensive and hard to replicate; yet, the essence of physics is repeatability. This work is an assessment of the Microscope results based on the laws of physics and knowledge from previous experiments, focusing on the limiting thermal noise and the treatment of acceleration outliers. Thermal noise reveals anomalies that we explain by stray sub-$\mu\rm V$ potentials caused by patch charges, giving rise to an unstable zero. The measurements were affected by numerous acceleration spikes occurring at the synodic frequencies relative to the Earth (the signal frequency) and the Sun, which we interpret as evidence of a thermal origin. In Microscope authors’ analysis, the spikes were removed and the resulting gaps replaced with artificial data (up to $35\%$, $40\%$ of the sessions data), which retain memory of the gaps and may simulate or cancel an effect (signal or systematic). An alternative approach based exclusively on real measured data would avoid any ambiguity. The lessons of Microscope are crucial to any future improved mission.
I Introduction
The Weak Equivalence Principle (WEP), or rather not a “principle” but the “hypothesis” of a complete physical equivalence of the gravitational field and the corresponding acceleration of the reference frame –as Einstein expressed it in 1907– is the founding pillar of General Relativity. In Newton’s formulation it was the equivalence of inertial to gravitational mass and the Universality of Free Fall (UFF).1 As such, it should be tested as precisely and accurately as possible, something physicists have been doing since Galileo’s time.
@@ I Introduction
The basic features of the Microscope experiment are addressed in Secs. II and III. Sec. II is concerned with the test masses, the accelerometers and the acceleration measurements, a potential violation signal from Earth, its signature and how it would be measured by a Sensor Unit (SU). In particular we show the $500$ gain factor in space and stress the difference between an inherently differential instrument like the torsion balance (no violation, no signal) and the SU of Microscope in which the acceleration difference is reconstructed from the individual measurements of two independent accelerometers. The relevance and consequences of such feature are addressed in Secs. III and IV.
Sec. III deals with the forces acting on the test masses, starting from drag and drag-free control and focusing on the electrostatic control of the test mass in the presence of Earth tidal forces, mechanical stiffness, biases, patch charges and fluctuating patch potentials.
Establishing the nature and source of the noise limiting the measurement is crucial for any experiment and a prerequisite for future improvements. In Sec. IV we analyze the limiting random noise. Since the preliminary results 5 the residual acceleration difference was reported to be limited by thermal noise with $1/\sqrt{\nu}$ frequency dependence attributed to internal damping in the gold wires. However, our analysis of the final results for all measurement sessions 4; 6 reveals anomalies which are inexplicable within the current best knowledge of this noise. In GOCE fluctuating patch potentials with $1/\sqrt{\nu}$ trend at low frequency were one of the main noise sources. We show that such a potential at sub-$\mu\rm V$ level at the frequency of the signal would be sufficient to explain the observed anomalies. This is because the fluctuating voltage multiplies a considerably large voltage needed to keep the test mass motionless at an arbitrary chosen zero of the sensor. As a result, a large force bias arises, hence, by way of the fluctuating patch potentials, a fluctuating zero. An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments.
@@
Systematic effects that were a matter of concern,7 turned out to be dominated by temperature variations. After calibration of the thermal response and a posteriori correction, residual systematic errors were much smaller than random errors or close to them.9; 8 This was achieved thanks to an effective multilayer thermal insulation blanket (MLI). However, MLI blankets are the likely source of large acceleration spikes (“glitches”) occurring simultaneously in the four test masses (Sec. V). GOCE was exempt thanks to an ad hoc, stiffer than usual, solution adopted for the insulation.
@@
An alternative approach relying solely on real measured data has been successfully used in the Eöt-Wash ground tests, with only $7\%$ of data removed.2 We argue that the same procedure is applicable to the Microscope test in orbit.
Conclusions are drawn in Sec. VI, along with the main takeaway lessons for future improvements.
II The accelerometers of Microscope and the signature of a potential WEP violation signal
@@
The concepts of equivalence between inertial and gravitational mass, UFF and WEP can all be used to mean the same thing.1 The theoretical framework is that of a 2-body problem (TM in the field of the source body) with WEP violation, and the violation signal lies in the orbit plane of the spacecraft.10 Since the sensor is rigid with the spacecraft it is natural to work in the non inertial reference frame of the spacecraft centered on its CoM relative to which –in case of space fixed attitude– the Earth orbits at angular velocity of modulus $\omega_{orb}=GM_{\oplus}/r^{3}$. Since the signal has two degrees of freedom it is desirable that the sensor too be sensitive in 2D (the orbit plane). The advantages over 1D sensors is to double the measurement data (and thus double the integration time $t_{int}$ and reduce the random noise by $\sqrt{2}$, since it is proportional to $1/\sqrt{t_{int}}$), and to avoid the risk of noise leakage from the other degree of freedom which is less sensitive.
@@
Two coaxial accelerometers form what is called a Sensor Unit (SU), and the spacecraft carries two of them (Fig. 1), one with different composition $\rm Pt$-$\rm Ti$ test cylinders for testing the WEP, named SUEP, and the other with $\rm Pt$-$\rm Pt$ cylinders named SUREF as a reference sensor because equal composition masses cannot violate the WEP even if their masses are different.
In each accelerometer two conductive electrode belts are placed at the two ends of the cylinder and measure the capacitance variations along the symmetry $X$ axis according to the “area-variation” scheme (Fig. 2).
@@
Like in all accelerometers by ONERA, the test masses of Microscope –although in the shape of hollow cylinders rather than cubes or parallelepipeds– are not allowed to move. A control voltage is applied to nullify any displacement so as to keep the TM motionless (except for the residual noise that cannot be controlled) at a chosen capacitive zero $O_{C}$ (which depends on geometry, construction, mounting and electronics). By calibration, a scale factor is established (in $\rm ms^{-2}/V$) by which the applied electric potential yields the acceleration acting on the test cylinder. Then, in Microscope a scale matching is performed between the two accelerometers of each SU and their acceleration difference is obtained, which is the physical quantity of interest for testing the WEP. This shows that the SU is not an intrinsically differential sensor, meaning that it does not measure the differential acceleration of the test masses directly, but provides it as the difference of two separate acceleration measurements (Secs. III and IV).
Figure 1: Microscope Sensor Units in space (left, figure taken from Microscope presentations) and WEP violation signal at zero spin (right). The orbit ($\simeq 700\,\rm km$ altitude) is near-circular ($e\simeq 10^{-4}$) and near-polar ($i\simeq 98.2^{\circ}$, making it sun-synchronous with solar panels facing the Sun, and the descending node at dawn). In the reference frame of the spacecraft, the $Z$ axis points at the CoM of the Earth and the $Y$ axis points in the direction of the outward orbit normal. The satellite spins clockwise (opposite to the orbital motion) so that the signal frequency is the sum of the orbit and spin rate. The separation between the centers of mass of the two Sensor Units is along the $Y$ axis, so that, projected on the $X,Z$ orbit plane, the position of the centers of mass of the test cylinders can be closer to the CoM of the satellite, as constrained by construction and mounting errors. At zero spin a WEP violation signal would act along the symmetry axis at $\omega_{orb}$ (right sketch, section of SUEP sensor on the orbit plane with the different composition cylinders depicted in different colors).
Figure 2: Microscope test cylinder with $X$ axis electrodes (Fig.5, Ref.(12)). In each accelerometer two conductive electrode belts are placed at the two ends of the cylinder with a gap of $598\,\mu\rm m$. If the CoM of the cylinder moves away from the central position, the area of the cylinder facing its electrode increases on one side and decreases on the other (“area-variation”), and so does the capacitance, by an amount depending on the displacement, hence on the acceleration that generated it.
@@
The advantage of rotation is to reduce the noise which is known to decrease at higher frequency (Sec. IV) and also to average thermal effects. However, as shown in both sketches of Fig. 1, by the very design of the Microscope sensor, rotation cannot occur around the symmetry axis which would ensure a higher spin rate and passive rotation by conservation of angular momentum, thus limiting attitude maneuvers only to keep the solar panels close to the perpendicular to the Sun.
III Forces acting on Microscope accelerometers and their control
@@
Ideally, being the same on all TMs (common mode effect) it should not affect the differential acceleration between the two TMs of the sensor unit. In reality, the two accelerometers of the SU do not respond with exactly the same acceleration. Only a limited rejection of common mode effects is achieved by scale matching in orbit. The residual differential effect of drag has the same frequency as the target signal and almost $90^{\circ}$ phase difference (at Microscope’s low altitude residual air dominates over solar radiation and the effect of drag is mostly along track, while a violation signal would point towards or away from Earth). A remaining differential component –after drag free control and partial rejection– with the same frequency and phase as the signal would mimic a violation signal; it is therefore necessary to ensure that the entire residual differential effect be smaller than the target.
@@
In most ONERA accelerometers (e.g. in GOCE) as well as in LISA, the TM is a cube (or a parallelepiped) and displacements are detected via capacitance variations in the classical “gap-variation” scheme in which the TM is located halfway between two capacitance plates facing its two opposite faces perpendicular to the sensitive axis. This scheme is used in Microscope for the $Y,Z$ axes, while the $X$ symmetry/sensitive axis is based on area-variation. As shown in Ref.(13), in gap-variation the electrostatic stiffness is negative and usually dominates over the mechanical, positive stiffness of the gold wire, hence the TM is unstable. Instead, in the area-variation scheme the electrostatic stiffness is smaller and positive, like the stiffness of the wire, and the TM is stable. In gap-variation the capacitance is inversely proportional to the gap squared, while in area-variation it is inversely proportional to the gap, but smaller gaps are known to give rise to higher disturbances. Instead of reducing the gap a larger surface of the electrodes is used to increase the capacity.
@@
The difference between $k_{\parallel}$ and $k_{\perp}$ is well illustrated in GOCE. In that case the TM of each accelerometer is a parallelepiped with two “ultra sensitive” axes ($Y,Z$) while the third one ($X$) is less sensitive. It is based on gap-variation and the gold wire ($5\,\rm\mu m$ diameter and $2.5\,\rm cm$ length) is directed along the $X$ axis. The stiffnesses reported are $k_{\parallel}=10^{-3}\,\rm N/m$ along $X$ and $k_{\perp}=1.8\times 10^{-6}\,\rm N/m$ along $Y,\,Z$; the gaps are of $32\,\rm\mu m$ along $X$ and $300\,\rm\mu m$ along $Y,\,Z$, and the larger face of the parallelepiped ($4\times 1\,\rm cm^{2}$) is perpendicular to the $X$ axis while the $Y,\,Z$ axes are perpendicular to the smaller faces (of $1\times 1\,\rm cm^{2}$).14 The $X$ axis was used to test the accelerometer in the lab, where the TM must be levitated against the local gravitational acceleration instead of the much smaller tidal acceleration in orbit.
The gold wire of Microscope’s cylinders has the same length and only a slightly bigger diameter than the wire of GOCE ($7\,\rm\mu m$ instead of $5$), hence the stiffness along the $X$ axis must be very close to $k_{\parallel}=10^{-3}\,\rm N/m$ reported for GOCE, and indeed the in orbit measurements found values around $\simeq 10^{-3}\,\rm N/m$ for all four test cylinders.12; 15
Instead,16 the stiffness expected for the Microscope wire was based on the theoretical formula:
@@
The reason for such a large discrepancy is that the theoretical stiffness of Eq.(2), as well as the experimental tests supporting it,17 apply to the response of the wire to a force acting perpendicularly to it, i.e. $k_{w}=k_{\perp}$, whereas the relevant stiffness is that acting along the symmetry/sensitive axis, $k_{\parallel}$.
Within the mission team, the in orbit measurements 15 have been questioned by Ref.(16), Sec. 5.1. However, the model used is incorrect, giving –for a wire along the $X$ axis– zero stiffness in the $Y,Z$ directions instead of the correct value resulting from Eq.(2).
A higher stiffness results in a higher level of random thermal noise from internal damping. It also affects the electrostatic control of the TM which keeps it motionless at a selected capacitive zero so that the voltage applied yields (through a scale factor in $\rm ms^{-2}/V$ established by calibration) the acceleration which caused the motion (Sec. IV).
@@
where $\omega_{orb}=\sqrt{GM_{\oplus}/r^{3}}=2\pi\nu_{orb}$ is the orbital angular velocity, $\vec{\varrho}=(X_{TM},\,Z_{TM})$ is the instantaneous position vector of the TM relative to the CoM of the spacecraft in the orbit plane and $\lambda=\omega_{orb}t$ is the instantaneous longitude relative to the spacecraft-to-Earth direction. The tidal acceleration is $\propto 1/r^{3}$ and acts at frequency $2\nu_{orb}$ while a violation signal (at zero spin) is at $\nu_{orb}$. If the orbit has a non zero eccentricity $10<e<$ 1, the effect is larger at pericenter than at apocenter and there is also a tidal component at $\nu_{orb}$, like the signal, although reduced by the factor $e$. If the spacecraft is spinning, tidal effects are up-converted to $2\nu_{{}_{EP}}$ and $\nu_{{}_{EP}}$.
@@
The difference of tidal accelerations between the two masses of each sensor unit is linear with the offsets $\Delta X,\,\Delta Z$ between their centers of mass within the SU. The largest effect at $2\nu_{EP}$ is then used to establish the values of these offsets, because tidal effects are deterministic and all the parameters they depend upon are measured, and this procedure allows a better determination than would be possible before launch. For SUEP the offsets measured in this way are $\Delta X\simeq 20\,\mu\rm m$ and $\Delta Z\simeq-5.6\,\mu\rm m$; for SUREF (in modulus) $\Delta Z$ is similar but $\Delta X$ is much larger ($\simeq-35.9\,\mu\rm m$).9 With SUEP measured values, the average tidal acceleration $\frac{1}{2}\omega_{orb}^{2}\Delta X$ as given in Eq.(3) is $\simeq 10^{-11}\,\rm ms^{-2}$.
@@
While the tidal effect at $2\nu_{EP}$ can be easily separated from the signal, the one at $\nu_{EP}$ competes with it. It is reduced by the very small orbital eccentricity of Microscope ($e\simeq 10^{-4}$). In addition, as mentioned in Ref.(18), it is almost cancelled when the spacecraft is spinning. This can be achieved by an appropriate choice of the spin frequency (it is an odd multiple of half the orbital frequency) and of the initial phase of the sensitive axis relative to the Earth so that, at pericenter and apocenter, when the tidal effect is maximum or minimum –resulting in the variation at $\nu_{{}_{EP}}$– it is $90^{\circ}$ away from the direction to the Earth, and therefore it is not affected by the tide (Fig. 1, right sketch). This effect is indeed shown negligible.19
@@
Under the tidal acceleration of Eq.(3) and the restoring elastic force due to $k_{\parallel}$ the TM will move, at any time, to an equilibrium position where the two effects balance each other, which occurs at an elongation $X_{eq}$ from the zero of the elastic force $0_{k}$. The electrostatic control acts to nullify this displacement and keep the TM at the selected capacitive zero $0_{C}$. If the two zeros do not coincide, the TM in $0_{C}$ is subjected to a constant elastic force that the control must nullify by applying a constant voltage (bias), in addition to the sinusoidal voltage required to nullify the tidal acceleration at $2\nu_{{}_{EP}}$.
@@
The resulting acceleration biases and their time evolution are reported, for all four accelerometers, in Ref.(15), Fig. 9. For SUEP the bias of the outer TM is $\simeq 10^{-6}\,\rm ms^{-2}$ and almost one order of magnitude smaller for the inner TM. With different offsets between their respective zeros, different masses and somewhat different stiffness, the acceleration bias is different for the two test masses of each sensor unit and the larger one dominates the acceleration difference. The time series for SUEP session #218 reported by Ref.(19), Slide 13, confirms that they are at the level of $10^{-6}\,\rm ms^{-2}$. For SUREF, the biases are of a few $10^{-7}\,\rm ms^{-2}$. 15

### v6

# Tel.: +39-050-2213254

@@ Abstract
The Weak Equivalence Principle is the founding pillar of General Relativity and as such it should be verified as precisely as possible. The Microscope experiment tested it in low-Earth orbit, finding that Pt and Ti test masses fall toward Earth with the same acceleration to about 1 part in $10^{15}$, an improvement of about two orders of magnitude over ground tests. Space missions, even if small, are expensive and hard to replicate; yet, the essence of physics is repeatability. This work is an assessment of the Microscope results based on the laws of physics and knowledge from previous experiments, focusing on the limiting thermal noise and the treatment of acceleration outliers. Thermal noise reveals anomalies that we explain by stray sub-$\mu\rm V$ potentials caused by patch charges, giving rise to an unstable zero. The measurements were affected by numerous acceleration spikes occurring at the synodic frequencies relative to the Earth (the signal frequency) and the Sun, which we interpret as evidence of a thermal origin. In Microscope authors’ analysis, the spikes were removed and the resulting gaps replaced with artificial data (up to $35\%$, $40\%$ of the sessions data), which retain memory of the gaps and may simulate or cancel an effect (signal or systematic). An alternative approach based exclusively on real measured data would avoid any ambiguity. The lessons of Microscope are crucial to any future improved mission.
Keywords:
General Relativity Equivalence Principle Precision Experiments in Space
pacs:
04.80Cc 07.87.+v
1 Introduction
The Weak Equivalence Principle (WEP), or rather not a “principle” but the “hypothesis” of a complete physical equivalence of the gravitational field and the corresponding acceleration of the reference frame –as Einstein expressed it in 1907– is the founding pillar of General Relativity. In Newton’s formulation it was the equivalence of inertial to gravitational mass and the Universality of Free Fall (UFF) (Nobili et al (2013)). As such, it should be tested as precisely and accurately as possible, something physicists have been doing since Galileo’s time.
@@ 1 Introduction
The basic features of the Microscope experiment are addressed in Secs. 2 and 3. Sec. 2 is concerned with the test masses, the accelerometers and the acceleration measurements, a potential violation signal from Earth, its signature and how it would be measured by a Sensor Unit (SU). In particular we show the $500$ gain factor in space and stress the difference between an inherently differential instrument like the torsion balance (no violation, no signal) and the SU of Microscope in which the acceleration difference is reconstructed from the individual measurements of two independent accelerometers. The relevance and consequences of such feature are addressed in Secs. 3 and 4.
Sec. 3 deals with the forces acting on the test masses, starting from drag and drag-free control and focusing on the electrostatic control of the test mass in the presence of Earth tidal forces, mechanical stiffness, biases, patch charges and fluctuating patch potentials.
Establishing the nature and source of the noise limiting the measurement is crucial for any experiment and a prerequisite for future improvements. In Sec. 4 we analyze the limiting random noise. Since the preliminary results (Touboul et al (2017)) the residual acceleration difference was reported to be limited by thermal noise with $1/\sqrt{\nu}$ frequency dependence attributed to internal damping in the gold wires. However, our analysis of the final results for all measurement sessions, which appeared at the same time in Touboul et al (2022a) and in a series of papers by the Microscope scientists on volume 39 of Classical and Quantum Gravity, reveals anomalies which are inexplicable within the current best knowledge of this noise. The GOCE mission (see Allasio et al (2010) for a general description) has shown that fluctuating patch potentials with $1/\sqrt{\nu}$ trend at low frequency were one of the main noise sources. We show that such a potential at sub-$\mu\rm V$ level at the frequency of the signal would be sufficient to explain the observed anomalies. This is because the fluctuating voltage multiplies a considerably large voltage needed to keep the test mass motionless at an arbitrary chosen zero of the sensor. As a result, a large force bias arises, hence, by way of the fluctuating patch potentials, a fluctuating zero. An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments.
@@
Systematic effects that were a matter of concern (Nobili and Anselmi (2018b)) turned out to be dominated by temperature variations. After calibration of the thermal response and a posteriori correction, residual systematic errors were much smaller than random errors or close to them (Touboul et al (2022b); Rodrigues et al (2022b)) This was achieved thanks to an effective multilayer thermal insulation blanket (MLI). However, MLI blankets are the likely source of large acceleration spikes (“glitches”) occurring simultaneously in the four test masses (Sec. 5). GOCE was exempt thanks to an ad hoc, stiffer than usual, solution adopted for the insulation.
@@
An alternative approach relying solely on real measured data has been successfully used in the Eöt-Wash ground tests, with only $7\%$ of data removed (Schlamminger et al (2008)). We argue that the same procedure is applicable to the Microscope test in orbit.
Conclusions are drawn in Sec. 6, along with the main takeaway lessons for future improvements.
2 The accelerometers of Microscope and the signature of a potential WEP violation signal
@@
The concepts of equivalence between inertial and gravitational mass, UFF and WEP can all be used to mean the same thing (Nobili et al (2013)). The theoretical framework is that of a 2-body problem (TM in the field of the source body) with WEP violation, and the violation signal lies in the orbit plane of the spacecraft (Nobili et al (2008)). Since the sensor is rigid with the spacecraft it is natural to work in the non inertial reference frame of the spacecraft centered on its CoM relative to which –in case of space fixed attitude– the Earth orbits at angular velocity of modulus $\omega_{orb}=\sqrt{GM_{\oplus}/r^{3}}$. Since the signal has two degrees of freedom it is desirable that the sensor too be sensitive in 2D (the orbit plane). The advantages over 1D sensors is to double the measurement data (and thus double the integration time $t_{int}$ and reduce the random noise by $\sqrt{2}$, since it is proportional to $1/\sqrt{t_{int}}$), and to avoid the risk of noise leakage from the other degree of freedom which in 1D sensors is less sensitive by design.
@@
Figure 1: The two Sensor Units of Microscope in orbit (left, figure taken from Microscope presentations) and the WEP violation signal at zero spin (right); figures not to scale. The orbit ($\simeq 700\,\rm km$ altitude) is near-circular ($e\simeq 10^{-4}$) and near-polar ($i\simeq 98.2^{\circ}$), making it sun-synchronous with solar panels facing the Sun and the descending node at dawn. In the reference frame of the spacecraft, $Y$ axis points in the direction of the outward orbit normal and the $X,Z$ axes lie in the orbit plane. The satellite spins clockwise (opposite to the orbital motion) so that the signal frequency is the sum of the orbit and spin rate. The separation between the centers of mass of the two Sensor Units is along the $Y$ axis, so that, projected on the $X,Z$ orbit plane, the position of the centers of mass of the test cylinders can be closer to the CoM of the satellite, as constrained by construction and mounting errors. At zero spin a WEP violation signal would act along the symmetry axis at $\omega_{orb}$ (right sketch, section of SUEP sensor on the orbit plane with the different composition cylinders depicted in different colors).
Figure 2: Microscope test cylinder with $X$ axis electrodes (Fig.5 in Touboul et al (2019)). In each accelerometer two conductive electrode belts are placed at the two ends of the cylinder with a gap of $598\,\mu\rm m$. If the CoM of the cylinder moves away from the central position, the area of the cylinder facing its electrode increases on one side and decreases on the other (“area-variation”), and so does the capacitance, by an amount depending on the displacement, hence on the acceleration that generated it.
@@
Two coaxial accelerometers form what is called a Sensor Unit (SU), and the spacecraft carries two of them (Fig. 1), one with different composition $\rm Pt$-$\rm Ti$ test cylinders for testing the WEP, named SUEP, and the other with $\rm Pt$-$\rm Pt$ cylinders named SUREF as a reference sensor, since WEP violation –if any– is expected to occur if the test bodies have different compositions (which may couple differently to the gravitational field of the source body), and not simply different masses.
In each accelerometer two conductive electrode belts are placed around the two ends of the cylinder and measure the capacitance variations along the symmetry $X$ axis according to the “area-variation” scheme (Fig. 2).
@@
Like in all accelerometers by ONERA, the test masses of Microscope –although in the shape of hollow cylinders rather than cubes or parallelepipeds– are not allowed to move. A control voltage is applied to nullify any displacement so as to keep the TM motionless (except for the residual noise that cannot be controlled) at a chosen capacitive zero $O_{C}$ (which depends on geometry, construction, mounting and electronics). By calibration, a scale factor is established (in $\rm ms^{-2}/V$) by which the applied electric potential yields the acceleration acting on the test cylinder. Then, in Microscope a scale matching is performed between the two accelerometers of each SU and their acceleration difference is obtained, which is the physical quantity of interest for testing the WEP. This shows that the SU is not an intrinsically differential sensor, meaning that it does not measure the differential acceleration of the test masses directly, but provides it as the difference of two separate acceleration measurements (Secs. 3 and 4).
@@
The advantage of rotation is to reduce the noise which is known to decrease at higher frequency (Sec. 4) and also to average thermal effects. However, as shown in both sketches of Fig. 1, by the very design of the Microscope sensor, rotation cannot occur around the symmetry axis, which would ensure a higher spin rate and passive rotation by conservation of angular momentum, in which case attitude manoeuvers are needed only to keep the solar panels close to the perpendicular to the Sun.
3 Forces acting on Microscope accelerometers and their control
@@
Ideally, being the same on all TMs (common mode effect) it should not affect the differential acceleration between the two TMs of the sensor unit. In reality, the two accelerometers of the SU do not respond with exactly the same acceleration. Only a limited rejection of common mode effects is achieved by scale matching in orbit. The residual differential effect of drag has the same frequency as the target signal and almost $90^{\circ}$ phase difference (at Microscope’s low altitude, residual air dominates over solar radiation and the effect of drag is mostly along track, while a violation signal would point towards or away from Earth). A remaining differential component –after drag free control and partial rejection– with the same frequency and phase as the signal would mimic a violation signal; it is therefore necessary to ensure that the entire residual differential effect be smaller than the target.
@@
In most ONERA accelerometers (e.g. in GOCE) as well as in LISA, the TM is a cube (or a parallelepiped) and displacements are detected via capacitance variations in the classical “gap-variation” scheme in which the TM is located halfway between two capacitance plates facing its two opposite faces perpendicular to the sensitive axis. This scheme is used in Microscope for the $Y,Z$ axes, while the $X$ symmetry/sensitive axis is based on area-variation. As shown in Josselin et al (1999), in gap-variation the electrostatic stiffness is negative and usually dominates over the mechanical, positive stiffness of the gold wire, hence the TM is unstable. Instead, in the area-variation scheme the electrostatic stiffness is much smaller than the stiffness of the wire and the TM is stable. In gap-variation the capacitance is inversely proportional to the gap squared, while in area-variation it is inversely proportional to the gap, but smaller gaps are known to give rise to higher disturbances. Instead of reducing the gap, a larger surface of the electrodes is used to increase the capacity.
@@
The difference between $k_{\parallel}$ and $k_{\perp}$ is well illustrated in GOCE. In that case the TM of each accelerometer is a parallelepiped with two “ultra sensitive” axes ($Y,Z$) while the third one ($X$) is less sensitive. It is based on gap-variation and the gold wire ($5\,\rm\mu m$ diameter and $2.5\,\rm cm$ length) is directed along the $X$ axis. The stiffnesses reported are $k_{\parallel}=10^{-3}\,\rm N/m$ along $X$ and $k_{\perp}=1.8\times 10^{-6}\,\rm N/m$ along $Y,\,Z$; the gaps are of $32\,\rm\mu m$ along $X$ and $300\,\rm\mu m$ along $Y,\,Z$, and the larger face of the parallelepiped ($4\times 1\,\rm cm^{2}$) is perpendicular to the $X$ axis while the $Y,\,Z$ axes are perpendicular to the smaller faces (of $1\times 1\,\rm cm^{2}$)(Catastini and Dumontel (2002)). The $X$ axis was used to test the accelerometer in the lab, where the TM must be levitated against the local gravitational acceleration instead of the much smaller tidal acceleration in orbit.
The gold wire of Microscope’s cylinders has the same length and only a slightly bigger diameter than the wire of GOCE ($7\,\rm\mu m$ instead of $5$), hence the stiffness along the $X$ axis must be very close to $k_{\parallel}=10^{-3}\,\rm N/m$ reported for GOCE, and indeed the in orbit measurements found values around $\simeq 10^{-3}\,\rm N/m$ for all four test cylinders (Touboul et al (2019); Chhun et al (2022)).
Instead, the stiffness expected for the Microscope wire (Bergé et al (2022b)) was based on the theoretical formula:
@@
The reason for such a large discrepancy is that the theoretical stiffness of Eq.(2), as well as the experimental tests supporting it (Willemenot and Touboul (2000)) apply to the response of the wire to a force acting perpendicularly to it, i.e. $k_{w}=k_{\perp}$, whereas the relevant stiffness is that acting along the symmetry/sensitive axis, $k_{\parallel}$.
Within the mission team, the in orbit measurements by Chhun et al (2022) have been questioned in Bergé et al (2022b), Sec. 5.1. However, the model used is incorrect, giving –for a wire along the $X$ axis– zero stiffness in the $Y,Z$ directions instead of the correct value resulting from Eq.(2).
A higher stiffness results in a higher level of random thermal noise from internal damping. It also affects the electrostatic control of the TM which keeps it motionless at a selected capacitive zero so that the voltage applied yields (through a scale factor in $\rm ms^{-2}/V$ established by calibration) the acceleration which caused the motion (Sec. 4).
@@
where $\omega_{orb}=\sqrt{GM_{\oplus}/r^{3}}=2\pi\nu_{orb}$ is the orbital angular velocity, $\lambda=\omega_{orb}t$ is the instantaneous longitude relative to the spacecraft-to-Earth direction and $\vec{\varrho}=(X_{TM},\,Z_{TM})$ is the instantaneous position vector of the TM relative to the CoM of the spacecraft in the orbit plane. If the orbit has a non zero eccentricity $10<e<$ 1, the effect is larger at pericenter than at apocenter and there is also a tidal component at $\nu_{orb}$, like the signal, although reduced by the factor $e$. If the spacecraft is spinning, tidal effects are up-converted to $2\nu_{{}_{EP}}$ and $\nu_{{}_{EP}}$.
@@
The difference of tidal accelerations between the two masses of each sensor unit is linear with the offsets $\Delta X,\,\Delta Z$ between their centers of mass within the SU. The largest effect at $2\nu_{EP}$ is then used to establish the values of these offsets, because tidal effects are deterministic and all the parameters they depend upon are measured, and this procedure allows a better determination than would be possible before launch. For SUEP the offsets measured in this way are $\Delta X\simeq 20\,\mu\rm m$ and $\Delta Z\simeq-5.6\,\mu\rm m$; for SUREF (in modulus) $\Delta Z$ is similar but $\Delta X$ is much larger ($\simeq-35.9\,\mu\rm m$) (Rodrigues et al (2022b)). With SUEP measured values, the average tidal acceleration $\frac{1}{2}\omega_{orb}^{2}\Delta X$ as given in Eq.(3) is $\simeq 10^{-11}\,\rm ms^{-2}$.
@@
While the tidal effect at $2\nu_{EP}$ can be easily separated from the signal, the one at $\nu_{EP}$ competes with it. It is reduced by the very small orbital eccentricity of Microscope ($e\simeq 10^{-4}$). In addition, as mentioned in Touboul et al (2012), it is almost cancelled when the spacecraft is spinning. This can be achieved by an appropriate choice of the spin frequency (it is an odd multiple of half the orbital frequency) and of the initial phase of the sensitive axis relative to the Earth so that, at pericenter and apocenter, when the tidal effect is maximum or minimum –resulting in the variation at $\nu_{{}_{EP}}$– it is $90^{\circ}$ away from the direction to the Earth, and therefore it is not affected by the tide (Fig. 1, right sketch). This effect is indeed shown negligible by M. Rodrigues in his talk at MGXV (15th Marcel Grossmann) meeting in Rome, 2018.
@@
Under the tidal acceleration of Eq.(3) and the restoring elastic force due to $k_{\parallel}$ the TM will move, at any time, to an equilibrium position where the two effects balance each other, which occurs at an elongation $X_{eq}$ from the zero of the elastic force $O_{k}$. The electrostatic control acts to nullify this displacement and keep the TM at the selected capacitive zero $O_{C}$. If the two zeros do not coincide, the TM in $O_{C}$ is subjected to a constant elastic force that the control must nullify by applying a constant voltage (bias), in addition to the sinusoidal voltage required to nullify the tidal acceleration at $2\nu_{{}_{EP}}$.
@@
The resulting acceleration biases and their time evolution are reported, for all four accelerometers, in Chhun et al (2022), Fig. 9. For SUEP the bias of the outer TM is $\simeq 10^{-6}\,\rm ms^{-2}$ and almost one order of magnitude smaller for the inner TM. With different offsets between their respective zeros, different masses and somewhat different stiffness, the acceleration bias is different for the two test masses of each sensor unit and the larger one dominates the acceleration difference. The time series for SUEP session #218 reported in Slide 13 of M. Rodrigues’ talk at MGXV in Rome, 2018, confirms that they are at the level of $10^{-6}\,\rm ms^{-2}$. For SUREF, the biases are of a few $10^{-7}\,\rm ms^{-2}$ (Chhun et al (2022)).